How much current must flow through a long straight wire for the magnetic field strength to be 1.0 mT at 1.0 cm from a wire? (?0 = 4?
Answer #1

B=1.0 mT=1.0*10-3 T

l= 1.0 cm =0.01 m

\mu _{0}=4\pi *10^{-7}

Lets I= current flow

B=\frac{\mu _{0}I}{2\pi l}

I=\frac{B*2\pi *l}{\mu _{0}}

I=\frac{1.0*10^{-3}*2\pi *1.0*10^{-2}m}{4\pi *10^{-7}}

I=50A

Answer is 50A


answered by: ANURANJAN SARSAM
Answer #2

B=mu*I/2*3.14*R

1*10-3=4*3.14*10-7*I/2*3.14*1*10-2

I=0.5*10-3/10-5

=50A